# Benchmarking Pretrained Molecular Embedding Models For Molecular Representation Learning

### v3

@@ Methods
We comprehensively evaluated molecular embeddings using 25 datasets - 7 from MoleculeNet (Wu et al. 2018) and 18 from TDC (Huang et al. 2021) - as they are among the most widely used benchmarks in molecular property prediction. Our focus is on classification tasks, as it is the most common approach in virtual screening (binary bioactivity classification) and ADMET prediction. Regression models were also very unstable in the initial experiments. The datasets span a broad range of tasks, including ADME, various toxicity endpoints, virtual screening, and more. In particular, TDC datasets include several small ones that reflect real-world ML-assisted drug discovery scenarios, where fine-tuning results in catastrophic overfitting. For a realistic out-of-distribution evaluation, we use a challenging scaffold-based train-test splits.
@@
We used the area under the receiver operating characteristic curve (AUROC) as an evaluation metric. AUROC is well-suited for imbalanced datasets, common in molecular ML, and is widely adopted in chemoinformatics, including in MoleculeNet and TDC benchmarks. The ECFP count fingerprint was used as the main baseline, due to its popularity, strong theoretical foundations, and consistently strong practical performance (Rogers and Hahn 2010). It also captures more information than its binary variant (Adamczyk and Ludynia 2024).
The benchmarking problem addressed in this work can be framed as a comparison of multiple models on multiple datasets (Demšar 2006). Bayesian testing is recommended for this setting (Benavoli et al. 2017), offering advantages over frequentist methods such as higher statistical power, interpretable posterior distributions, and the ability to define a region of practical equivalence (ROPE). Specifically, we use the hierarchical Bayesian Bradley-Terry (BBT) model (Wainer 2023), which is designed for multimodel comparisons. It provides clear decisions on whether one model outperforms another, is practically equivalent, or if the evidence is inconclusive, while offering strong theoretical grounding and intuitive probabilistic interpretation. We follow the recommendations of (Wainer 2023) for model interpretation, but expand the ROPE to 25257575%. That is, if the probability that one model outperforms another in a given dataset falls within $[0.25,0.75]$, we consider them practically equivalent. This makes the comparisons more conservative, reducing false positives when evaluating many models. Despite this much stricter criterion, the BBT model still yields very confident rankings, as detailed in the Results section.
@@
Figure 1 illustrates our approach. This setup enables the rigorous evaluation of numerous models on a large-scale, diverse benchmark under a fair comparison framework, with analysis conducted using the dedicated BBT model. See the Supplementary Material for details on datasets, model variants, hyperparameter grids, hardware, and further explanation of Bayesian tests and the BBT model.
Results
@@
The aggregated results of the model, the average rank, and the AUROC, are shown in Table 1. The key finding is that only four models outperformed the ECFP fingerprint. The top-performing model, CLAMP, is a fusion of molecular fingerprints. The other three - R-MAT, MolBERT, and ChemBERTa - use distinct architectures and pretraining strategies. We note that the MTR variant, incorporating more chemistry-specific knowledge into pretraining, vastly outperformed the MLM variant in all cases (see the Supplementary Material for details), and thus it is also used here.
@@
ChemBERTa
No decision
can be made
CDDD
@@
Bayesian testing
Aggregated metrics alone, such as mean AUROC and mean rank, are insufficient for a fair comparison of the models (Demšar 2006; Wainer 2023). Mean AUROC is sensitive to outliers (e.g., datasets unusually easy for certain models), while mean rank ignores the magnitude of performance differences. To address these limitations, we use the BBT model, designed specifically for robust multi-model comparisons. The results, with ECFP as the baseline, are shown in Table 2. Following the procedure described in (Wainer 2023), we categorize models as: better than ECFP, equivalent (within the ROPE), worse or undecidable (insufficient evidence). All posterior probabilities are provided in the Supplementary Material.
These results show that only the CLAMP model is statistically significantly better than ECFP. Further analyses in the following sections confirm that CLAMP is the only model that consistently outperforms, or at least matches, other approaches.
ChemBERTa is practically equivalent to ECFP, but comes with significantly higher computational and memory costs. For CDDD, the evidence is inconclusive, meaning it performs much better or worse than ECFP depending on the dataset. A key practical advantage of hierarchical Bayesian testing is its ability to capture such nuanced relationships while accounting for multiple hypothesis testing, something that traditional null hypothesis testing cannot do (Benavoli et al. 2017).
Despite having a higher average rank and AUROC, R-MAT and MolBERT are deemed worse than ECFP by the BBT model, suggesting that their apparent performance may be driven by outliers on just a few datasets. R-MAT was closer to equivalence, with a probability 80% of falling within ROPE. All other models, including AP and TT molecular fingerprints, are significantly worse than ECFP. This does not mean that these models always perform poorly, but that their overall likelihood of outperforming ECFP is low. Such insights help guide model selection, reducing engineering effort and computational cost.
We also note that we used a conservative ROPE range of 2525-7575%. Despite this wide interval, the BBT model remains very confident in its conclusions, often assigning a probability near 100%. The detailed tables are available in the Supplementary Material.
Inter-model win rates
Figure 2 provides a detailed analysis of win rates between models, showing the percentage of datasets where one model outperforms another. Models are ranked by their number of wins against the ECFP baseline. To exclude insignificant differences, following (Benavoli et al. 2017), we treat AUROC differences below 0.01% as ties (practical equivalence). We also apply this for all subsequent analyses.
@@
Per-dataset wins against ECFP
@@
Discussion and practical remarks
@@
Qualitatively, reproducibility and code quality strongly correlated with model performance. The best models featured easy-to-run, well-documented code and dependencies. Molecular fingerprint baselines implemented with scikit-fingerprints (Adamczyk and Ludynia 2024) were also very user-friendly. In contrast, models requiring substantial engineering effort to run, such as GraphFP, GEM, or GraphMVP, often delivered disappointing results.
@@
    We echo existing literature (Benavoli et al. 2017; Wainer 2023) recommending Bayesian statistical tests, particularly the Bayesian Bradley-Terry (BBT) model, for multi-model, multi-dataset comparisons, as they provide clear aggregation and intuitive interpretation of results.
Conclusions
We presented a comprehensive benchmark of pretrained molecular representation models. Analysis of 25 models on 25 datasets, based on a fair comparison setting and supported with a Bayesian Bradley-Terry model, challenges perceived progress in this domain and indicates inadequate evaluation rigor in previous works. All models, except CLAMP, are either worse than or practically equivalent to the baseline ECFP fingerprint on molecular property prediction tasks. Furthermore, the representations learned by GNNs are of significantly worse quality than those of other approaches, including molecular fingerprints. Those findings call for further in-depth empirical and theoretical analyses, to increase the practical relevance of ML models for drug development and chemoinformatics.
Supplementary Material
Source code
The source code for the benchmarking procedure and the implementation of the referenced models is available on GitHub at https://github.com/scikit-fingerprints/benchmarking_molecular_models. The repository also contains AUROC scores for each embedding model, dataset, and classification head.
Datasets
@@ Conclusions
Dataset Benchmark No. samples No. tasks Mean share of positive samples
@@
Classifiers and hyperparameter tuning
@@
Hyperparameter grids were designed to result in reasonable computational time, and also give good results based on initial experiments. For Random Forest, we tune min_samples_split in range $[2,4,6,8,10$, and always use 500 trees with entropy loss function, following e.g. Local Topological Profile (Adamczyk and Czech 2023) and MOLTOP (Adamczyk and Czech 2024). For logistic regression, we tune reverse regularization strength $\lambda$, as 10 values on logarithmic scale from $10^{-2}$ to $10^{3}$. For kNN, we tune number of neighbors k in range $[1,3,5,7,9]$.
Hardware used
@@
Bayesian statistical tests
@@
Bayesian Bradley–Terry model
Bayesian procedures return the posterior probability that one model is better than another and allow explicit probability statements about practical equivalence, overcoming well-known limitations of frequentist NHST in this setting, e.g., p-values confound effect size and sample size; non-significance does not imply equivalence (Benavoli et al. 2017). Let $P(i\succ j)$ be a Bayesian posterior-predictive probability: “the probability that model $i$ beats model jj”. $P(i\succ j)$ is the central quantity reported by the Bayesian Bradley–Terry model (Wainer 2023): it combines the empirical evidence (observed wins), model assumptions (the logistic link), and parameter uncertainty into a single, decision-relevant probability.
Pairwise win counts
For every ordered pair of models $(i,j)$ let $W_{ij}$ be the (possibly fractional) [^1] Ties are handled by the “spread” rule; each tie contributes $0.5$ to both $W_{ij}$ and $W_{ji}$. number of datasets on which model $i$ outperforms model $j$. We denote the total number of comparisons by:
$N_{ij}\;=\;W_{ij}+W_{ji}.$
Likelihood
The win counts are modeled with a binomial Bradley–Terry likelihood:
$W_{ij}\;\sim\;\mathrm{Binomial}\!\Bigl{(}N_{ij},\,\pi_{ij}\Bigr{)},$ (1)
Priors
Latent abilities $\beta_{i}$ are given a hierarchical Gaussian prior with shared scale $\sigma$:
$\beta_{i}\;\sim\;\mathcal{N}\!\bigl{(}0,\sigma^{2}\bigr{)},$ (3)
$\sigma\;\sim\;\mathrm{LogNormal}\!\bigl{(}0,0.5^{2}\bigr{)},$ (4)
$\sum_{i=1}^{M}\beta_{i}\;=\;0\quad\text{(identifiability)}.$ (5)
Posterior inference
We sample from the joint posterior $p\!\bigl{(}\boldsymbol{\beta},\sigma\mid\{W_{ij}\}\bigr{)}$ by Markov-Chain Monte Carlo (Stan implementation). Convergence is checked, and posterior predictive checks (PPCs) are used to confirm that the observed win counts are plausible draws from the fitted model.
Posterior summaries
For each ordered pair $(i,j)$ and each MCMC draw $s$,
$\pi_{ij}^{(s)}\;=\;\operatorname{logit}^{-1}\!\bigl{(}\beta_{i}^{(s)}-\beta_{j}^{(s)}\bigr{)}.$
we report: (i) the posterior mean $\hat{\pi}_{ij}$, (ii) the $89$ % highest-density interval (HDI), (iii) $P_{\text{in~ROPE}}$, the posterior mass lying inside the prespecified probability ROPE $\rho_{L},\rho_{U}]$, (iv) $P_{\text{above~0.5}}$, the probability that model $i$ truly outperforms model $j$. Global abilities are summarised by the posterior means $\hat{\beta}_{i}=\mathbb{E}[\beta_{i}]$, which induce a unique ranking, while retaining full pairwise uncertainty.
Decision rule
Let $\rho_{L},\rho_{U}]\!=\![0.25,0.75]$ be the probability-space region of practical equivalence (ROPE):
  • $\text{declare }i\text{ better than }j\;\Longleftarrow\;\hat{\pi}_{ij}>\rho_{U},$
  • $\text{declare equivalence}\;\Longleftarrow\;P_{\text{in~ROPE}}\geq 0.95$.
Pairs that satisfy neither condition are deemed inconclusive.
@@
Decision rules and ROPE
The BBT supports a universal probability ROPE, independent of the downstream metric’s scale. Wainer proposes $[0.45,0.55]$: if the posterior lies largely in $[0.45,0.55]$ models $i$ and $j$ are practically equivalent. Researchers may widen or narrow this interval depending on conservativeness. Following Wainer’s guidance for interpretation but opting for stronger conservativeness when many models are compared, we used a wide probability ROPE of $[0.25,0.75]$. Despite this strict criterion, posterior probabilities were often near $1.0$ or $0.0$, yielding confident decisions.
Step-by-step protocol
  1. 1.
  2. 2.
    Within-dataset pairwise comparisons: For each data set and each pair of models $(i,j)$: if $|AUROC_{i}-AUROC_{j}|<\epsilon$ ($\epsilon=0.01$, i.e. 1%) record a tie; else, record a win for the higher AUROC model.
  3. 3.
  4. 4.
    Fit the BBT: Use the binomial/logistic likelihood with hierarchical priors $\beta_{i}\sim\mathcal{N}\!\bigl{(}0,\sigma^{2}\bigr{)}$,
    $\sigma\sim\mathrm{LogNorm}\!\bigl{(}0,0.5^{2}\bigr{)}$; sample via MCMC to obtain posterior draws of $\{\beta_{i}\}$. Check convergence and perform PPC.
  5. 5.
    Summarize pairwise posteriors: For every pair: $(i,j)$ transform draws of $\beta_{i}-\beta_{j}$ to draws of $P(i\succ j)$, report the posterior mean, $89\textendash 95$% highest density intervals (HDIs), $P_{\text{in~ROPE}}$, and $P_{\text{above~0.5}}$.
  6. 6.
    Make decisions with ROPE:
    • If the posterior of $P(i\succ j)$ lies predominantly within $[0.25,0.75]$ declare $i$ and $j$ practically equivalent (our conservative setting)
    • Else if mean $P(i\succ j)>0.75$ declare $i$ better than $j$; if <0.25<0.25, declare worse.
  7. 7.
    Aggregate and rank: Rank models by posterior mean $\beta_{i}$ (or by their expected win probability against the field) and report pairwise probabilities alongside the ranking to convey uncertainty.

### v4

@@ 3 Methods
We comprehensively evaluated molecular embeddings using 25 datasets - 7 from MoleculeNet MoleculeNet and 18 from TDC TDC - as they are among the most widely used benchmarks in molecular property prediction. Our focus is on classification tasks, as it is the most common approach in virtual screening (binary bioactivity classification) and ADMET prediction. This also allows direct comparison of models using the same metric and easy application of Bayesian statistical tests. The datasets span a broad range of tasks, including ADME, various toxicity endpoints, virtual screening, and more. In particular, TDC datasets include several small ones that reflect real-world ML-assisted drug discovery scenarios, where fine-tuning results in catastrophic overfitting. For a realistic out-of-distribution evaluation, we use a challenging scaffold-based train-test splits.
@@
We used the area under the receiver operating characteristic curve (AUROC) as an evaluation metric. AUROC is well-suited for imbalanced datasets, common in molecular ML, and is widely adopted in chemoinformatics, including in MoleculeNet and TDC benchmarks. The ECFP count fingerprint was used as the main baseline, due to its popularity, strong theoretical foundations, and consistently strong practical performance fp_ecfp . It also captures more information than its binary variant scikit-fingerprints .
@@
Figure 1 illustrates our approach. This setup enables the rigorous evaluation of numerous models on a large-scale, diverse benchmark under a fair comparison framework, with analysis conducted using the dedicated Bayesian statistical testing, based on Bayesian Bradley-Terry model. See the Supplementary Material for details on datasets, model variants, hyperparameter grids, and hardware.
3.1 Bayesian Bradley-Terry model
The benchmarking problem addressed in this work can be framed as a comparison of multiple models on multiple datasets Demsar . Aggregated metrics alone, such as mean AUROC and mean rank, are insufficient for a fair comparison of the models under such setting Demsar ; BBT . Mean AUROC is sensitive to outliers (e.g., datasets unusually easy for certain models), while mean rank ignores the magnitude of performance differences. Instead, Bayesian testing is recommended for this setting bayesian_tests , offering advantages over frequentist methods such as higher statistical power, interpretable posterior distributions, and the ability to define a region of practical equivalence (ROPE). Bayesian procedures return the posterior probability that one model is better than another and allow explicit probability statements about practical equivalence. They also overcome the well-known limitations of frequentist null hypothesis statistical tests (NHST), such as p-values confounding effect size and sample size, or the fact that non-significance for NHST does not imply model equivalence bayesian_tests . Instead, Bayesian tests model $P(i\succ j)$, the posterior-predictive probability: “the probability that model $i$ beats model jj”. Further, the notion of ROPE allows returning a tie (practical equivalence) as the result of statistical test, both on a given dataset and in the final ranking.
Specifically, we use the hierarchical Bayesian Bradley-Terry (BBT) model BBT , which is designed for robust multimodel comparisons on multiple datasets. It provides clear decisions on whether one model outperforms another, is practically equivalent, or if the evidence is inconclusive, while offering strong theoretical grounding and intuitive probabilistic interpretation. This statistical testing model combines empirical evidence (observed wins), model assumptions (logistic link), and parameter uncertainty into a single probability relevant to decision. $P(i\succ j)$ is the central quantity reported by this model.
Below, we summarize the BBT model, and refer the interested readers to bayesian_tests and BBT for detailed descriptions and proofs of statistical properties.
Pairwise win counts
For every ordered pair of models $(i,j)$ let $W_{ij}$ be the number of datasets on which model $i$ outperforms model $j$ in a given metric, e.g. AUROC. If this difference is too small, e.g. less than $1\%$, models are deemed practically equivalent, adding $0.5$ to both $W_{ij}$ and $W_{ji}$ (known as spread rule for ties). We denote the total number of comparisons by $N_{ij}\;=\;W_{ij}+W_{ji}$.
Likelihood
The win counts are modeled with a binomial Bradley–Terry likelihood:
$W_{ij}\;\sim\;\mathrm{Binomial}\!\Bigl(N_{ij},\,\pi_{ij}\Bigr),$ (1)
Priors
Latent abilities $\beta_{i}$ of each model, i.e. their true general performance, are given a hierarchical Gaussian prior with shared scale $\sigma$:
$\beta_{i}\;\sim\;\mathcal{N}\!\bigl(0,\sigma^{2}\bigr),$ (3)
$\sigma\;\sim\;\mathrm{LogNormal}\!\bigl(0,0.5^{2}\bigr),$ (4)
$\sum_{i=1}^{M}\beta_{i}\;=\;0\quad\text{(identifiability)}.$ (5)
Posterior inference
To obtain the posterior distribution of probability that model $i$ is better than model $j$, we employ the Markov-Chain Monte Carlo approach. Specifically, the joint posterior is $p\bigl(\boldsymbol{\beta},\sigma\mid\{W_{ij}\}\bigr)$. Convergence is checked and posterior predictive checks (PPCs) are used to confirm that the observed win counts are plausible draws from the fitted model.
Posterior summaries
For each ordered pair $(i,j)$ and each MCMC draw $s$,
$\pi_{ij}^{(s)}\;=\;\operatorname{logit}^{-1}\!\bigl(\beta_{i}^{(s)}-\beta_{j}^{(s)}\bigr)$
we report:
  1. 1.
    the posterior mean $\hat{\pi}_{ij}$,
  2. 2.
    $P_{\text{in~ROPE}}$, the posterior mass lying within the prespecified probability ROPE $\rho_{L},\rho_{U}]$,
  3. 3.
    $P_{\text{above~0.5}}$, the probability that model $i$ truly outperforms model $j$.
Global abilities are summarized by the posterior means $\hat{\beta}_{i}=\mathbb{E}[\beta_{i}]$, which induce a unique ranking while retaining full pairwise uncertainty.
Decision rule and model ranking
Let $\rho_{L},\rho_{U}]$ be the probability-space region of practical equivalence (ROPE). This measure is directly measured on the probability distribution and independent of the scale of the model quality metric (e.g. AUROC). Depending on values of $\rho_{L}$ and $\rho_{U}$, we control the conservativeness of the model, with wider ROPE interval making the model more probable to declare inconclusive evidence based on the data provided.
The BBT makes the decision about the relative quality of the two models as follows:
  • if $P_{\text{in~ROPE}}\geq 0.95$, models are equivalent
  • if $\hat{\pi}_{ij}>\rho_{U}$, then model $i$ is better than $j$
  • if $\hat{\pi}_{ij}<\rho_{L}$, then model $j$ is better than $i$
  • otherwise, evidence is inconclusive, and no decision can be made
In addition to the pairwise model comparisons, BBT allows ordering the models by their posterior mean of $\beta_{i}$. This yields a unique global ranking while retaining pairwise posteriors for uncertainty-aware claims.
3.2 Step-by-step protocol
Here, we summarize the procedure of Bayesian statistical testing with BBT for our benchmarking protocol.
We follow the recommendations of BBT for model interpretation, but expand the ROPE to 35356565%. That is, if the probability that one model outperforms another in a given dataset falls within $[0.35,0.65]$, we consider them practically equivalent. This reduces false positives when evaluating many models. Despite this much stricter criterion, the BBT model still yields very confident rankings, with most probabilities close to $100\%$.
  1. 1.
  2. 2.
    Pairwise comparisons within the data set: For each dataset and each pair of models $(i,j)$, declare a win for model with higher AUROC (or a tie if the difference is $1\%$).
  3. 3.
  4. 4.
    Fit the BBT: Use the binomial/logistic likelihood with hierarchical priors $\beta_{i}\sim\mathcal{N}\bigl(0,\sigma^{2}\bigr)$, $\sigma\sim\mathrm{LogNorm}\!\bigl(0,0.5^{2}\bigr)$, and sample via MCMC to obtain posterior draws of $\{\beta_{i}\}$. Check convergence and perform PPC.
  5. 5.
    Summarize pairwise posteriors: For every pair $(i,j)$, transform draws of $\beta_{i}-\beta_{j}$ to draws of $P(i\succ j)$, report the posterior mean, $P_{\text{in~ROPE}}$, and $P_{\text{above~0.5}}$.
  6. 6.
    Make decisions with ROPE: if $P_{\text{in~ROPE}}\geq 0.95$, declare models $i$ and $j$ equivalent. If $\hat{\pi}_{ij}>0.65$, declare $i$ better if $\hat{\pi}_{ij}<0.35$, declare $j$ better. Otherwise, declare inconclusive evidence for those models.
  7. 7.
    Aggregate and rank: Classify models by posterior mean $\beta_{i}$ (their expected win probability) and report pairwise probabilities alongside the ranking to convey uncertainty.
4 Results
@@
The aggregated results of the model, the average rank, and the AUROC, are shown in Table 1. The key finding is that only four models outperformed the ECFP fingerprint. The top-performing model, CLAMP, is a fusion of molecular fingerprints. The other three - R-MAT, MolBERT, and ChemBERTa - use distinct architectures and pretraining strategies. We note that the MTR variant, incorporating more chemistry-specific knowledge in the pretraining, vastly outperformed the MLM variant in all cases (see the Supplementary Material for details), and thus it is also used here.
@@
R-MAT, MolBERT, ChemBERTa,
CDDD, Atom Pair, MAT
@@
4.1 Bayesian testing results
As discussed in the Methods section, comparison mean AUROC or ranks is insufficient for proper comparison of multiple models on multiple datasets. Thus, we employ the BBT model, and report its results in Table 2, with ECFP as the baseline. Following the procedure described in BBT , we categorize models as: better than ECFP, equivalent (within the ROPE), worse or undecidable (insufficient evidence). All posterior probabilities are provided in the Supplementary Material.
These results show that only the CLAMP model is statistically significantly better than ECFP. Further analyzes in the following sections confirm that CLAMP is the only model that consistently outperforms, or at least matches, other approaches.
Several transformer-based approaches including R-MAT, MolBERT, ChemBERTa and MAT are practically equivalent to ECFP, but coming with significantly higher computational cost. A key practical advantage of hierarchical Bayesian testing is its ability to capture such nuanced relationships while adjusting for multiple hypothesis testing, something traditional null hypothesis testing cannot do bayesian_tests . Additionally, RNN-based CDDD model and Atom Pair fingerprint also performed on par with ECFP.
All other models, including TT molecular fingerprints, are significantly worse than ECFP. This does not mean that these models always perform poorly, but that their overall probability of outperforming ECFP is low. Such insights help guide model selection, reducing engineering effort and computational cost.
We also note that we used a conservative ROPE range of 3535-6565%. Despite this wide interval, the BBT model remains very confident in its conclusions, often assigning a probability near 100%. The detailed tables and experiments with other ROPE intervals are available in the Supplementary Material.
4.2 Inter-model win rates
Figure 2 provides a detailed analysis of win rates between models, showing the percentage of datasets where one model outperforms another. Models are ranked by their number of wins against the ECFP baseline. To exclude insignificant differences, following bayesian_tests , we treat AUROC differences below 0.01% as ties (practical equivalence). We also apply this for all subsequent analyses.
@@ 4.2 Inter-model win rates
4.3 Per-dataset wins against ECFP
@@
4.4 Discussion and practical remarks
@@
Qualitatively, reproducibility and code quality strongly correlated with model performance. The best models featured easy-to-run, well-documented code and dependencies. Molecular fingerprint baselines implemented with scikit-fingerprints scikit-fingerprints were also very user-friendly. In contrast, models requiring substantial engineering effort to run, such as GraphFP, GEM, or GraphMVP, often delivered disappointing results.
@@
    We echo existing literature bayesian_tests ; BBT recommending Bayesian statistical tests, particularly the Bayesian Bradley-Terry (BBT) model, for multi-model, multi-dataset comparisons, as they provide clear aggregation and intuitive interpretation of results.
5 Conclusions
We presented a comprehensive benchmark of pretrained molecular representation models. Analysis of 25 models on 25 datasets, based on a fair comparison setting and supported with a Bayesian Bradley-Terry model, challenges perceived progress in this domain and indicates inadequate evaluation rigor in previous works. All models, except CLAMP, are either worse than or practically equivalent to the baseline ECFP fingerprint on molecular property prediction tasks. Furthermore, the representations learned by GNNs are of significantly worse quality than those of other approaches, including molecular fingerprints. Those findings call for further in-depth empirical and theoretical analyzes, to increase the practical relevance of ML models for drug development and chemoinformatics.
6 Acknowledgements
We thank BIT Student Scientific Group for computational resources and aid for this project. We thank Michał Lampert for initial experiments and analyses. We further thank Aleksandra Elbakyan for her work and support for accessibility of science.
6.1 Data availability
All datasets and benchmarks used are publicly available at https://github.com/scikit-fingerprints/benchmarking_molecular_models.
7 Author contributions statement
Contributions follow the CRediT format. M.P. was responsible for software, validation, data curation, and writing. J.A. was responsible for conceptualization, methodology, investigation, data curation, and writing.W.C. was responsible for writing, review, and supervision.
8 Competing interests
No competing interest is declared.
9 Funding
This work was supported by the general funding from the Ministry of Science and Higher Education to AGH University in Krakow, and by project Excellence Initiative – Research University (IDUB) for AGH University of Krakow.
Supplementary information
10 Datasets
@@
Dataset Benchmark No. samples No. tasks
Mean share
of positive samples
@@ 11 Model variants
12 Classifiers and hyperparameter tuning
@@
Hyperparameter grids were designed to result in reasonable computational time, and also give good results based on initial experiments. For Random Forest, we tune min_samples_split in range $[2,4,6,8,10$, and always use 500 trees with entropy loss function, following e.g. Local Topological Profile [1] and MOLTOP [2]. For logistic regression, we tune reverse regularization strength $\lambda$, as 10 values on logarithmic scale from $10^{-2}$ to $10^{3}$. For kNN, we tune number of neighbors k in range $[1,3,5,7,9]$.
13 Hardware used
@@
14 Bayesian statistical tests
@@ 14.1 Data preparation
14.2 Decision rules and ROPE
The BBT supports a universal probability ROPE, independent of the downstream metric’s scale. Wainer proposes a default interval of $[0.45,0.55]$: if the posterior lies largely within this region, models (i) and (j) are considered practically equivalent. Researchers may widen or narrow this interval depending on the desired level of conservativeness. In our study, we followed Wainer’s guidance but opted for a more conservative ROPE of $[0.35,0.65]$ due to the large number of models under comparison. Despite this strict threshold, posterior probabilities were often close to $1.0$ or $0.0$, allowing confident decisions.
Table 6 reports the full BBT statistics comparing all models against ECFP, including the posterior mean, mean difference (delta), probability of exceeding $0.50$, and probability of lying within the chosen ROPE. To illustrate how conclusions depend on the ROPE definition, Table 5 presents model interpretations across a range of ROPE intervals, from $[0.45,0.55]$ to $[0.20,0.80]$.
ROPE Better than ECFP Practically equivalent Unknown Worse than ECFP
(0.45, 0.55) CLAMP, RMAT None MolBERT, CDDD, ChemBERTa, MAT, AtomPair Rest
(0.4, 0.6) CLAMP, RMAT MolBERT, CDDD, ChemBERTa, MAT AtomPair Rest
(0.35, 0.65) CLAMP RMAT, MolBERT, CDDD, ChemBERTa, MAT, AtomPair None Rest
(0.3, 0.7) None CLAMP, RMAT, MolBERT, CDDD, ChemBERTa, MAT, AtomPair, MoLFormer, Mol2Vec None Rest
(0.25, 0.75) None CLAMP, RMAT, MolBERT, CDDD, ChemBERTa, MAT, AtomPair, MoLFormer, Mol2Vec, TT None Rest
(0.2, 0.8) None CLAMP, RMAT, MolBERT, CDDD, ChemBERTa, MAT, AtomPair, MoLFormer, Mol2Vec, TT, ChemFM, Uni-Mol2 None Rest
Table 5: BBT model interpretations for different ROPE values.